If a sum of money becomes 4 times more over a period of 10 years, what is the rate of simple interest?

-30%

-20%

-40%

-Not enough information to solve problem
Answer #1

Answer : 30 %

Explanation :

Let the Principle be $100 & rate of simple interest be X %.

Now

T = 10 years

P = $100

A = $400

Therefore Interest = A - P = 400 - 100 = $300.

Now we know that simple interest = (P * R * T) / 100

or, 300 = (100 * X * 10 ) / 100

or X = 30 %